xx CottonmouthBSSN

CottonmouthBSSN

Lucas Sanches Max Morris

Steven R. Brandt

Date

Abstract

Provides the BSSN equations as generated by the EinsteinEngine.

1 Introduction

The Einstein Engine is a DSL and toolkit for creating Cactus thorns. It can generate complete CarpetX thorns from simple recipes written in Python.

1.1 How to set up EinsteinEngine Create a venv and activate it.

xxxpythonx-mxvenvxvenvx&&x.x./venv/bin/activate
xxx

Install the dependencies.

xxxpythonx-mxpipxinstallx-rxrequirements.txt
xxx

Install EinsteinEngine.

xxxpythonx-mxpipxinstallx.
xxx

1.2 How to generate a thorn

Be sure the venv is activated.

xxx.x./venv/bin/activate
xxx

Typecheck the recipe with MyPy.

xxxmypyxrecipes/Cottonmouth/Z4c.py
xxx

Run the recipe.

     xxxpythonxrecipes/Cottonmouth/Z4c.pyx[--vacuum]x[--fd-orderxN]
     xxx

2 Parameters

By default, the code is generated with matter terms on. To disable them, run using the –vacuum flag.

By default, the code will be generated with 4th order finite difference stencils and 5th order Kreiss-Oliger dissipation. Use the –fd-order flag to switch this to Nth order finite difference stencils and (N+1)th Kreiss-Oliger dissipation.

3 Parameters

11 applyxnewradx 10 Scope: restricted01  BOOLEAN

11 Description: Apply NewRadX boundary conditions

11                  10                01  Default: false

11

11 chixfloor  10 Scope: restricted01  REAL

11 Description: Chi will never be smaller than this value

11  Range  10                01  Default: 1e-06

11 *  10

11

11 dissipationxepsilon10 Scope: restricted01  REAL

11 Description: The ammount of dissipation to add.

11  Range  10                01  Default: 0.32

11 *  10

11

11 etaxbeta  10 Scope: restricted01  REAL

11 Description: Standard Gamma driver eta gauge parameter. Must be of order 2 / MxADM

11  Range  10                01  Default: 2.0

11 *  10

11 evolvedxlapsexfloor10 Scope: restricted01  REAL

11 Description: The evolved lapse will never be smaller than this value

11  Range  10                01  Default: 1e-08

11 *  10

11

11 kappax1  10 Scope: restricted01  REAL

11 Description: Constraint damping parameter kappax1. Must be of order 1 / L wehre L is the typical simulation scale.

11  Range  10                01  Default: 0.02

11 *  10

11

11 kappax2  10 Scope: restricted01  REAL

11 Description: Constraint damping parameter kappax2.

11  Range  10                01  Default: 0.0

11 *  10

11

11 radpowerxat  10 Scope: restricted01  REAL

11 Description: NewRadX radpower for Atxij

11  Range  10                01  Default: 1.0

11 *  10

11

11 radpowerxchi  10 Scope: restricted01  REAL

11 Description: NewRadX radpower for chi

11  Range  10                01  Default: 1.0

11 *  10

11

11 radpowerxevoxgammat 10 Scope: restricted01  REAL

11 Description: NewRadX radpower Gammax

11  Range  10                01  Default: 1.0

11 *  10

11

11 radpowerxevoxlapse10 Scope: restricted01  REAL

11 Description: NewRadX radpower for the lapse, alpha

11  Range  10                01  Default: 1.0

11 *  10

11

11 radpowerxevoxshift10 Scope: restricted01  REAL

11 Description: NewRadX radpower for the shift, betax

11  Range  10                01  Default: 1.0

11 *  10

11

11 radpowerxgt  10 Scope: restricted01  REAL

11 Description: NewRadX radpower for gtxij

11  Range  10                01  Default: 1.0

11 *  10

11

11 radpowerxtheta 10 Scope: restricted01  REAL

11 Description: NewRadX radpower for Theta

11  Range  10                01  Default: 1.0

11 *  10

11

11 radpowerxtrk  10 Scope: restricted01  REAL

11 Description: NewRadX radpower the trace of Kxij

11  Range  10                01  Default: 1.0

11 *  10

4 Interfaces

General Implements: cottonmouthz4c4m

Inherits: admbasex driver odesolvers tmunubasex

Grid Variables

4.0.1 PUBLIC GROUPS

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11  at  10                    01  centering 11 centering={0 0 0}
11                    10  AtDD00  01  compact 11 0
11                    10  AtDD01  01 dimensions 11 3
11                    10  AtDD02  01 distribution 11 DEFAULT
11                    10  AtDD11  01 group type 11 GF
11                    10  AtDD12  01  tags 11 rhs=”CottonmouthZ4c4m::Atxrhs” parities={+1 +1 +1 -1 -1 +1 -1 +1 -1 +1 +1 +1 +1 -1 -1 +1 +1 +1}
11                    10  AtDD22  01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  atxrhs  10                    01  centering 11 centering={0 0 0}
11                    10  AtxrhsDD00  01  compact 11 0
11                    10  AtxrhsDD01  01 dimensions 11 3
11                    10  AtxrhsDD02  01 distribution 11 DEFAULT
11                    10  AtxrhsDD11  01 group type 11 GF
11                    10  AtxrhsDD12  01  tags 11 checkpoint=”no” parities={+1 +1 +1 -1 -1 +1 -1 +1 -1 +1 +1 +1 +1 -1 -1 +1 +1 +1}
11                    10  AtxrhsDD22  01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  hamcons  10                    01  centering 11 centering={0 0 0}
11                    10  HamCons  01  compact 11 0
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 checkpoint=”no” parities={+1 +1 +1}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  momcons  10                    01  centering 11 centering={0 0 0}
11                    10  MomConsU0  01  compact 11 0
11                    10  MomConsU1  01 dimensions 11 3
11                    10  MomConsU2  01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 checkpoint=”no” parities={-1 +1 +1 +1 -1 +1 +1 +1 -1}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  theta  10                    01  centering 11 centering={0 0 0}
11                    10  Theta  01  compact 11 0
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 rhs=”CottonmouthZ4c4m::Thetaxrhs” parities={+1 +1 +1}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  thetaxrhs  10                    01  centering 11 centering={0 0 0}
11                    10  Thetaxrhs  01  compact 11 0
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 checkpoint=”no” parities={+1 +1 +1}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11  ztcons  10                    01  centering 11 centering={0 0 0}
11                    10  ZtConsU0  01  compact 11 0
11                    10  ZtConsU1  01 dimensions 11 3
11                    10  ZtConsU2  01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 checkpoint=”no” parities={-1 +1 +1 +1 -1 +1 +1 +1 -1}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  chi  10                    01  centering 11 centering={0 0 0}
11                    10  chi  01  compact 11 0
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 rhs=”CottonmouthZ4c4m::chixrhs” parities={+1 +1 +1}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  chixrhs  10                    01  centering 11 centering={0 0 0}
11                    10  chixrhs  01  compact 11 0
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 checkpoint=”no” parities={+1 +1 +1}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 evoxgammat 10 01 centering 11 centering={0 0 0}
11                    10 evoxGammatU0 01  compact 11 0
11                    10 evoxGammatU1 01 dimensions 11 3
11                    10 evoxGammatU2 01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 rhs=”CottonmouthZ4c4m::evoxGammatxrhs” parities={-1 +1 +1 +1 -1 +1 +1 +1 -1}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 evoxgammatxrhs 10                    01  centering 11 centering={0 0 0}
11                    10 evoxGammatxrhsU0 01  compact 11 0
11                    10 evoxGammatxrhsU1 01 dimensions 11 3
11                    10 evoxGammatxrhsU2 01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 checkpoint=”no” parities={-1 +1 +1 +1 -1 +1 +1 +1 -1}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  evoxlapse  10                    01  centering 11 centering={0 0 0}
11                    10  evoxlapse  01  compact 11 0
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 rhs=”CottonmouthZ4c4m::evoxlapsexrhs” parities={+1 +1 +1}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11  evoxlapsexrhs  10                    01  centering 11 centering={0 0 0}
11                    10  evoxlapsexrhs 01  compact 11 0
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 checkpoint=”no” parities={+1 +1 +1}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  evoxshift  10                    01  centering 11 centering={0 0 0}
11                    10  evoxshiftU0  01  compact 11 0
11                    10  evoxshiftU1  01 dimensions 11 3
11                    10  evoxshiftU2  01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 rhs=”CottonmouthZ4c4m::evoxshiftxrhs” parities={-1 +1 +1 +1 -1 +1 +1 +1 -1}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11 evoxshiftxrhs 10 01 centering 11 centering={0 0 0}
11                    10 evoxshiftxrhsU0 01  compact 11 0
11                    10 evoxshiftxrhsU1 01 dimensions 11 3
11                    10 evoxshiftxrhsU2 01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 checkpoint=”no” parities={-1 +1 +1 +1 -1 +1 +1 +1 -1}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  gt  10                    01  centering 11 centering={0 0 0}
11                    10  gtDD00  01  compact 11 0
11                    10  gtDD01  01 dimensions 11 3
11                    10  gtDD02  01 distribution 11 DEFAULT
11                    10  gtDD11  01 group type 11 GF
11                    10  gtDD12  01  tags 11 rhs=”CottonmouthZ4c4m::gtxrhs” parities={+1 +1 +1 -1 -1 +1 -1 +1 -1 +1 +1 +1 +1 -1 -1 +1 +1 +1}
11                    10  gtDD22  01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  gtxrhs  10                    01  centering 11 centering={0 0 0}
11                    10  gtxrhsDD00  01  compact 11 0
11                    10  gtxrhsDD01  01 dimensions 11 3
11                    10  gtxrhsDD02  01 distribution 11 DEFAULT
11                    10  gtxrhsDD11  01 group type 11 GF
11                    10  gtxrhsDD12  01  tags 11 checkpoint=”no” parities={+1 +1 +1 -1 -1 +1 -1 +1 -1 +1 +1 +1 +1 -1 -1 +1 +1 +1}
11                    10  gtxrhsDD22  01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  trk  10                    01  centering 11 centering={0 0 0}
11                    10  trK  01  compact 11 0
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 rhs=”CottonmouthZ4c4m::trKxrhs” parities={+1 +1 +1}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL

11

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11  trkxrhs  10                    01  centering 11 centering={0 0 0}
11                    10  trKxrhs  01  compact 11 0
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01  tags 11 checkpoint=”no” parities={+1 +1 +1}
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL

11

Uses header: newradx.hxx 5 Schedule This section lists all the variables which are assigned storage by thorn Cottonmouth/CottonmouthZ4c4m. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11 Always:  11 x
11 At[1]  11 x
11 Atxrhs[1]  11 x
11 evoxGammatxrhs[1] 11 x
11 evoxlapse[1]  11 x
11 evoxlapsexrhs[1]  11 x
11 evoxshift[1]  11 x
11 evoxshiftxrhs[1]  11 x
11 gt[1]  11 x
11 gtxrhs[1]  11 x
11 trK[1]  11 x
11 trKxrhs[1]  11 x
11 HamCons[1]  11 x
11 MomCons[1]  11 x
11 Theta[1]  11 x
11 Thetaxrhs[1]  11 x
11 ZtCons[1]  11 x
11 chi[1]  11 x
11 chixrhs[1]  11 x
11 evoxGammat[1]  11 x

11 x  11 x

11

Scheduled Functions

CottonmouthZ4cxInitialGroup x admxtoxz4cxpt1xgroup x group for function ‘admxtoxz4cxpt1‘. generated by einsteinengine. x

11 x 11 Language: 11 c
11 x 11 Type:  11 group

11

admxtoxz4cxpt1xgroup x admxtoxz4cxpt1 x function ‘admxtoxz4cxpt1‘ generated by einsteinengine.

11 x 11 Language: 11 c
11 x 11 Reads:  11 admbasex::alp(interior)
11 x 11 x  11 admbasex::betax(interior)
11 x 11 x  11 admbasex::betay(interior)
11 x 11 x  11 admbasex::betaz(interior)
11 x 11 x  11 admbasex::gxx(interior)
11 x 11 x  11 admbasex::gxy(interior)
11 x 11 x  11 admbasex::gxz(interior)
11 x 11 x  11 admbasex::gyy(interior)
11 x 11 x  11 admbasex::gyz(interior)
11 x 11 x  11 admbasex::gzz(interior)
11 x 11 x  11 admbasex::kxx(interior)
11 x 11 x  11 admbasex::kxy(interior)
11 x 11 x  11 admbasex::kxz(interior)
11 x 11 x  11 admbasex::kyy(interior)
11 x 11 x  11 admbasex::kyz(interior)
11 x 11 x  11 admbasex::kzz(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 atdd00(interior)
11 x 11 x  11 atdd01(interior)
11 x 11 x  11 atdd02(interior)
11 x 11 x  11 atdd11(interior)
11 x 11 x  11 atdd12(interior)
11 x 11 x  11 atdd22(interior)
11 x 11 x  11 theta(interior)
11 x 11 x  11 chi(interior)
11 x 11 x  11 evoxlapse(interior)
11 x 11 x  11 evoxshiftu0(interior)
11 x 11 x  11 evoxshiftu1(interior)
11 x 11 x  11 evoxshiftu2(interior)
11 x 11 x  11 gtdd00(interior)
11 x 11 x  11 gtdd01(interior)
11 x 11 x  11 gtdd02(interior)
11 x 11 x  11 gtdd11(interior)
11 x 11 x  11 gtdd12(interior)
11 x 11 x  11 gtdd22(interior)
11 x 11 x  11 trk(interior)

11

CottonmouthZ4cxPostStepGroup x z4cxenforcexpt2xgroup x group for function ‘z4cxenforcexpt2‘. generated by einsteinengine. x

11 x 11 After:  11 z4cxenforcexpt1xgroup
11 x 11 Language: 11 c
11 x 11 Type:  11 group

11

z4cxenforcexpt2xgroup x z4cxenforcexpt2 x function ‘z4cxenforcexpt2‘ generated by einsteinengine.

11 x 11 Language: 11 c
11 x 11 Reads:  11 atdd00(interior)
11 x 11 x  11 atdd01(interior)
11 x 11 x  11 atdd02(interior)
11 x 11 x  11 atdd11(interior)
11 x 11 x  11 atdd12(interior)
11 x 11 x  11 atdd22(interior)
11 x 11 x  11 gtdd00(interior)
11 x 11 x  11 gtdd01(interior)
11 x 11 x  11 gtdd02(interior)
11 x 11 x  11 gtdd11(interior)
11 x 11 x  11 gtdd12(interior)
11 x 11 x  11 gtdd22(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 atdd00(interior)
11 x 11 x  11 atdd01(interior)
11 x 11 x  11 atdd02(interior)
11 x 11 x  11 atdd11(interior)
11 x 11 x  11 atdd12(interior)
11 x 11 x  11 atdd22(interior)

11

CottonmouthZ4cxRHSGroup x z4cxrhsxgroup x group for function ‘z4cxrhs‘. generated by einsteinengine. x

11 x 11 Language: 11 c
11 x 11 Type:  11 group

11

z4cxrhsxgroup x z4cxrhs x function ‘z4cxrhs‘ generated by einsteinengine.

11 x 11 Language: 11 c
11 x 11 Reads:  11 atdd00(everywhere)
11 x 11 x  11 atdd01(everywhere)
11 x 11 x  11 atdd02(everywhere)
11 x 11 x  11 atdd11(everywhere)
11 x 11 x  11 atdd12(everywhere)
11 x 11 x  11 atdd22(everywhere)
11 x 11 x  11 theta(everywhere)
11 x 11 x  11 tmunubasex::ettt(interior)
11 x 11 x  11 tmunubasex::ettx(interior)
11 x 11 x  11 tmunubasex::etty(interior)
11 x 11 x  11 tmunubasex::ettz(interior)
11 x 11 x  11 tmunubasex::etxx(interior)
11 x 11 x  11 tmunubasex::etxy(interior)
11 x 11 x  11 tmunubasex::etxz(interior)
11 x 11 x  11 tmunubasex::etyy(interior)
11 x 11 x  11 tmunubasex::etyz(interior)
11 x 11 x  11 tmunubasex::etzz(interior)
11 x 11 x  11 chi(everywhere)
11 x 11 x  11 evoxgammatu0(everywhere)
11 x 11 x  11 evoxgammatu1(everywhere)
11 x 11 x  11 evoxgammatu2(everywhere)
11 x 11 x  11 evoxlapse(everywhere)
11 x 11 x  11 evoxshiftu0(everywhere)
11 x 11 x  11 evoxshiftu1(everywhere)
11 x 11 x  11 evoxshiftu2(everywhere)
11 x 11 x  11 gtdd00(everywhere)
11 x 11 x  11 gtdd01(everywhere)
11 x 11 x  11 gtdd02(everywhere)
11 x 11 x  11 gtdd11(everywhere)
11 x 11 x  11 gtdd12(everywhere)
11 x 11 x  11 gtdd22(everywhere)
11 x 11 x  11 trk(everywhere)
11 x 11 Type:  11 function
11 x 11 Writes: 11 atxrhsdd00(interior)
11 x 11 x  11 atxrhsdd01(interior)
11 x 11 x  11 atxrhsdd02(interior)
11 x 11 x  11 atxrhsdd11(interior)
11 x 11 x  11 atxrhsdd12(interior)
11 x 11 x  11 atxrhsdd22(interior)
11 x 11 x  11 thetaxrhs(interior)
11 x 11 x  11 chixrhs(interior)
11 x 11 x  11 evoxgammatxrhsu0(interior)
11 x 11 x  11 evoxgammatxrhsu1(interior)
11 x 11 x  11 evoxgammatxrhsu2(interior)
11 x 11 x  11 evoxlapsexrhs(interior)
11 x 11 x  11 evoxshiftxrhsu0(interior)
11 x 11 x  11 evoxshiftxrhsu1(interior)
11 x 11 x  11 evoxshiftxrhsu2(interior)
11 x 11 x  11 gtxrhsdd00(interior)
11 x 11 x  11 gtxrhsdd01(interior)
11 x 11 x  11 gtxrhsdd02(interior)
11 x 11 x  11 gtxrhsdd11(interior)
11 x 11 x  11 gtxrhsdd12(interior)
11 x 11 x  11 gtxrhsdd22(interior)
11 x 11 x  11 trkxrhs(interior)

11

CottonmouthZ4cxPostStepGroup x z4cxtoxadmxgroup x group for function ‘z4cxtoxadm‘. generated by einsteinengine.

11 x 11 After:  11 z4cxenforcexpt2xgroup
11 x 11 Language: 11 c
11 x 11 Type:  11 group

11

z4cxtoxadmxgroup x z4cxtoxadm x function ‘z4cxtoxadm‘ generated by einsteinengine. x

11 x 11 Language: 11 c
11 x 11 Reads:  11 atdd00(everywhere)
11 x 11 x  11 atdd01(everywhere)
11 x 11 x  11 atdd02(everywhere)
11 x 11 x  11 atdd11(everywhere)
11 x 11 x  11 atdd12(everywhere)
11 x 11 x  11 atdd22(everywhere)
11 x 11 x  11 chi(everywhere)
11 x 11 x  11 evoxlapse(everywhere)
11 x 11 x  11 evoxshiftu0(everywhere)
11 x 11 x  11 evoxshiftu1(everywhere)
11 x 11 x  11 evoxshiftu2(everywhere)
11 x 11 x  11 gtdd00(everywhere)
11 x 11 x  11 gtdd01(everywhere)
11 x 11 x  11 gtdd02(everywhere)
11 x 11 x  11 gtdd11(everywhere)
11 x 11 x  11 gtdd12(everywhere)
11 x 11 x  11 gtdd22(everywhere)
11 x 11 x  11 trk(everywhere)
11 x 11 Type:  11 function
11 x 11 Writes: 11 admbasex::alp(everywhere)
11 x 11 x  11 admbasex::betax(everywhere)
11 x 11 x  11 admbasex::betay(everywhere)
11 x 11 x  11 admbasex::betaz(everywhere)
11 x 11 x  11 admbasex::gxx(everywhere)
11 x 11 x  11 admbasex::gxy(everywhere)
11 x 11 x  11 admbasex::gxz(everywhere)
11 x 11 x  11 admbasex::gyy(everywhere)
11 x 11 x  11 admbasex::gyz(everywhere)
11 x 11 x  11 admbasex::gzz(everywhere)
11 x 11 x  11 admbasex::kxx(everywhere)
11 x 11 x  11 admbasex::kxy(everywhere)
11 x 11 x  11 admbasex::kxz(everywhere)
11 x 11 x  11 admbasex::kyy(everywhere)
11 x 11 x  11 admbasex::kyz(everywhere)
11 x 11 x  11 admbasex::kzz(everywhere)

11

ODESolversxInitial x cottonmouthz4cxinitialgroup x z4 initialization routines x

11 x 11 After: 11 admbasexxpostinitial
11 x 11 Type: 11 group

11

ODESolversxPostStep x cottonmouthz4cxpoststepgroup x z4 post-step routines x

11 x 11 Before: 11 admbasexxsetadmvars
11 x 11 Type: 11 group

ODESolversxRHS x cottonmouthz4cxrhsgroup x z4 equations rhs computation x

11 x 11 Type: 11 group

11

CCTKxANALYSIS x cottonmouthz4cxanalysisgroup x z4 analysis routines x

11 x 11 Type: 11 group

11

CottonmouthZ4cxInitialGroup (conditional) x admxtoxz4cxpt2xgroup x group for function ‘admxtoxz4cxpt2‘. generated by einsteinengine. x

11 x 11 After:  11 admxtoxz4cxpt1xgroup
11 x 11 Language: 11 c
11 x 11 Type:  11 group

11

ODESolversxPostStep x syncxstate x empty function for explicitly syncing variables as provided in an explicitsyncbatch. syncs: at, theta, chi, evoxgammat, evoxlapse, evoxshift, gt, trk. generated by einsteinengine. x

11 x 11 Before: 11 admbasexxsetadmvars
11 x 11 Language: 11 c
11 x 11 Sync:  11 at
11 x 11 x  11 theta

11 x 11 x  11 chi

11 x 11 x  11 evoxgammat
11 x 11 x  11 evoxlapse
11 x 11 x  11 evoxshift

11 x 11 x  11 gt
11 x 11 x  11 trk

11 x 11 Type:  11 function

11

z4cxtoxadmxgroup x syncxz4c x empty function for explicitly syncing variables as provided in an explicitsyncbatch. syncs: at, gt, chi, evoxlapse, evoxshift, trk. generated by einsteinengine. x

11 x 11 Before: 11 z4cxtoxadm
11 x 11 Language: 11 c
11 x 11 Sync:  11 at

11 x 11 x  11 gt
11 x 11 x  11 chi

11 x 11 x  11 evoxlapse
11 x 11 x  11 evoxshift

11 x 11 x  11 trk

11 x 11 Type:  11 function

11

admxtoxz4cxpt2xgroup x syncxz4cxpt2 x empty function for explicitly syncing variables as provided in an explicitsyncbatch. syncs: gt. generated by einsteinengine.

11 x 11 Before: 11 admxtoxz4cxpt2
11 x 11 Language: 11 c
11 x 11 Sync:  11 gt
11 x 11 Type:  11 function

11

CottonmouthZ4cxRHSGroup (conditional) x z4cxapplyxnewradxxtheta x function for applying newradx boundary conditions: theta -> thetaxrhs. generated by einsteinengine. x

11 x 11 After:  11 z4cxapplyxdissipationxgroup
11 x 11 Language: 11 c
11 x 11 Reads:  11 theta(interior)
11 x 11 x  11 thetaxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 thetaxrhs(interior)

11

CottonmouthZ4cxRHSGroup (conditional) x z4cxapplyxnewradxxchi x function for applying newradx boundary conditions: chi -> chixrhs. generated by einsteinengine. x

11 x 11 After:  11 z4cxapplyxdissipationxgroup
11 x 11 Language: 11 c
11 x 11 Reads:  11 chi(interior)
11 x 11 x  11 chixrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 chixrhs(interior)

11

CottonmouthZ4cxRHSGroup (conditional) x z4cxapplyxnewradxxtrk x function for applying newradx boundary conditions: trk -> trkxrhs. generated by einsteinengine. x

11 x 11 After:  11 z4cxapplyxdissipationxgroup
11 x 11 Language: 11 c
11 x 11 Reads:  11 trk(interior)
11 x 11 x  11 trkxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 trkxrhs(interior)

11

CottonmouthZ4cxRHSGroup (conditional) x z4cxapplyxnewradxxevoxgammat x function for applying newradx boundary conditions: evoxgammat -> evoxgammatxrhs. generated by einsteinengine. x

11 x 11 After:  11 z4cxapplyxdissipationxgroup
11 x 11 Language: 11 c
11 x 11 Reads:  11 evoxgammat(interior)
11 x 11 x  11 evoxgammatxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 evoxgammatxrhs(interior)

11

CottonmouthZ4cxRHSGroup (conditional) x z4cxapplyxnewradxxgtxxx x function for applying newradx boundary conditions: gt[l0, l0] -> gtxrhs[l0, l0]. generated by einsteinengine.

11 x 11 After:  11 z4cxapplyxdissipationxgroup
11 x 11 Language: 11 c
11 x 11 Reads:  11 gt(interior)
11 x 11 x  11 gtxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 gtxrhs(interior)

11

CottonmouthZ4cxRHSGroup (conditional) x z4cxapplyxnewradxxgtxxy x function for applying newradx boundary conditions: gt[l0, l1] -> gtxrhs[l0, l1]. generated by einsteinengine. x

11 x 11 After:  11 z4cxapplyxdissipationxgroup
11 x 11 Language: 11 c
11 x 11 Reads:  11 gt(interior)
11 x 11 x  11 gtxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 gtxrhs(interior)

11

CottonmouthZ4cxRHSGroup (conditional) x z4cxapplyxnewradxxgtxxz x function for applying newradx boundary conditions: gt[l0, l2] -> gtxrhs[l0, l2]. generated by einsteinengine. x

11 x 11 After:  11 z4cxapplyxdissipationxgroup
11 x 11 Language: 11 c
11 x 11 Reads:  11 gt(interior)
11 x 11 x  11 gtxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 gtxrhs(interior)

11

admxtoxz4cxpt2xgroup (conditional) x admxtoxz4cxpt2 x function ‘admxtoxz4cxpt2‘ generated by einsteinengine. x

11 x 11 Language: 11 c
11 x 11 Reads:  11 gtdd00(everywhere)
11 x 11 x  11 gtdd01(everywhere)
11 x 11 x  11 gtdd02(everywhere)
11 x 11 x  11 gtdd11(everywhere)
11 x 11 x  11 gtdd12(everywhere)
11 x 11 x  11 gtdd22(everywhere)
11 x 11 Type:  11 function
11 x 11 Writes: 11 evoxgammatu0(interior)
11 x 11 x  11 evoxgammatu1(interior)
11 x 11 x  11 evoxgammatu2(interior)

11

CottonmouthZ4cxRHSGroup (conditional) x z4cxapplyxnewradxxgtxyy x function for applying newradx boundary conditions: gt[l1, l1] -> gtxrhs[l1, l1]. generated by einsteinengine.

11 x 11 After:  11 z4cxapplyxdissipationxgroup
11 x 11 Language: 11 c
11 x 11 Reads:  11 gt(interior)
11 x 11 x  11 gtxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 gtxrhs(interior)

11

CottonmouthZ4cxRHSGroup (conditional) x z4cxapplyxnewradxxgtxyz x function for applying newradx boundary conditions: gt[l1, l2] -> gtxrhs[l1, l2]. generated by einsteinengine. x

11 x 11 After:  11 z4cxapplyxdissipationxgroup
11 x 11 Language: 11 c
11 x 11 Reads:  11 gt(interior)
11 x 11 x  11 gtxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 gtxrhs(interior)

11

CottonmouthZ4cxRHSGroup (conditional) x z4cxapplyxnewradxxgtxzz x function for applying newradx boundary conditions: gt[l2, l2] -> gtxrhs[l2, l2]. generated by einsteinengine. x

11 x 11 After:  11 z4cxapplyxdissipationxgroup
11 x 11 Language: 11 c
11 x 11 Reads:  11 gt(interior)
11 x 11 x  11 gtxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 gtxrhs(interior)

11

CottonmouthZ4cxRHSGroup (conditional) x z4cxapplyxnewradxxat x function for applying newradx boundary conditions: at -> atxrhs. generated by einsteinengine. x

11 x 11 After:  11 z4cxapplyxdissipationxgroup
11 x 11 Language: 11 c
11 x 11 Reads:  11 at(interior)
11 x 11 x  11 atxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 atxrhs(interior)

11

CottonmouthZ4cxRHSGroup (conditional) x z4cxapplyxnewradxxevoxlapse x function for applying newradx boundary conditions: evoxlapse -> evoxlapsexrhs. generated by einsteinengine. x

11 x 11 After:  11 z4cxapplyxdissipationxgroup
11 x 11 Language: 11 c
11 x 11 Reads:  11 evoxlapse(interior)
11 x 11 x  11 evoxlapsexrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 evoxlapsexrhs(interior)

11

CottonmouthZ4cxRHSGroup (conditional) x z4cxapplyxnewradxxevoxshift x function for applying newradx boundary conditions: evoxshift -> evoxshiftxrhs. generated by einsteinengine. x

11 x 11 After:  11 z4cxapplyxdissipationxgroup
11 x 11 Language: 11 c
11 x 11 Reads:  11 evoxshift(interior)
11 x 11 x  11 evoxshiftxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 evoxshiftxrhs(interior)

11

CCTKxPOSTPOSTINITIAL x initxevolvexcottonmouthz4c4m x group containing all functions from evolve that should run before step 0. generated by einsteinengine. x

11 x 11 Language: 11 c
11 x 11 Type:  11 group

11

CCTKxPOSTPOSTINITIAL x initxanalysisxcottonmouthz4c4m x group containing all functions from analysis that should run before step 0. generated by einsteinengine. x

11 x 11 After:  11 initxevolvexcottonmouthz4c4m
11 x 11 Language: 11 c
11 x 11 Type:  11 group

11

CottonmouthZ4cxRHSGroup x z4cxapplyxdissipationxgroup x group for function ‘z4cxapplyxdissipation‘. generated by einsteinengine. x

11 x 11 After:  11 z4cxrhsxgroup
11 x 11 Language: 11 c
11 x 11 Type:  11 group

11

z4cxapplyxdissipationxgroup x z4cxapplyxdissipation x function ‘z4cxapplyxdissipation‘ generated by einsteinengine.

11 x 11 Language: 11 c
11 x 11 Reads:  11 atdd00(everywhere)
11 x 11 x  11 atdd01(everywhere)
11 x 11 x  11 atdd02(everywhere)
11 x 11 x  11 atdd11(everywhere)
11 x 11 x  11 atdd12(everywhere)
11 x 11 x  11 atdd22(everywhere)
11 x 11 x  11 atxrhsdd00(interior)
11 x 11 x  11 atxrhsdd01(interior)
11 x 11 x  11 atxrhsdd02(interior)
11 x 11 x  11 atxrhsdd11(interior)
11 x 11 x  11 atxrhsdd12(interior)
11 x 11 x  11 atxrhsdd22(interior)
11 x 11 x  11 theta(everywhere)
11 x 11 x  11 thetaxrhs(interior)
11 x 11 x  11 chi(everywhere)
11 x 11 x  11 chixrhs(interior)
11 x 11 x  11 evoxgammatu0(everywhere)
11 x 11 x  11 evoxgammatu1(everywhere)
11 x 11 x  11 evoxgammatu2(everywhere)
11 x 11 x  11 evoxgammatxrhsu0(interior)
11 x 11 x  11 evoxgammatxrhsu1(interior)
11 x 11 x  11 evoxgammatxrhsu2(interior)
11 x 11 x  11 evoxlapse(everywhere)
11 x 11 x  11 evoxlapsexrhs(interior)
11 x 11 x  11 evoxshiftu0(everywhere)
11 x 11 x  11 evoxshiftu1(everywhere)
11 x 11 x  11 evoxshiftu2(everywhere)
11 x 11 x  11 evoxshiftxrhsu0(interior)
11 x 11 x  11 evoxshiftxrhsu1(interior)
11 x 11 x  11 evoxshiftxrhsu2(interior)
11 x 11 x  11 gtdd00(everywhere)
11 x 11 x  11 gtdd01(everywhere)
11 x 11 x  11 gtdd02(everywhere)
11 x 11 x  11 gtdd11(everywhere)
11 x 11 x  11 gtdd12(everywhere)
11 x 11 x  11 gtdd22(everywhere)
11 x 11 x  11 gtxrhsdd00(interior)
11 x 11 x  11 gtxrhsdd01(interior)
11 x 11 x  11 gtxrhsdd02(interior)
11 x 11 x  11 gtxrhsdd11(interior)
11 x 11 x  11 gtxrhsdd12(interior)
11 x 11 x  11 gtxrhsdd22(interior)
11 x 11 x  11 trk(everywhere)
11 x 11 x  11 trkxrhs(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 atxrhsdd00(interior)
11 x 11 x  11 atxrhsdd01(interior)
11 x 11 x  11 atxrhsdd02(interior)
11 x 11 x  11 atxrhsdd11(interior)
11 x 11 x  11 atxrhsdd12(interior)
11 x 11 x  11 atxrhsdd22(interior)
11 x 11 x  11 thetaxrhs(interior)
11 x 11 x  11 chixrhs(interior)
11 x 11 x  11 evoxgammatxrhsu0(interior)
11 x 11 x  11 evoxgammatxrhsu1(interior)
11 x 11 x  11 evoxgammatxrhsu2(interior)
11 x 11 x  11 evoxlapsexrhs(interior)
11 x 11 x  11 evoxshiftxrhsu0(interior)
11 x 11 x  11 evoxshiftxrhsu1(interior)
11 x 11 x  11 evoxshiftxrhsu2(interior)
11 x 11 x  11 gtxrhsdd00(interior)
11 x 11 x  11 gtxrhsdd01(interior)
11 x 11 x  11 gtxrhsdd02(interior)
11 x 11 x  11 gtxrhsdd11(interior)
11 x 11 x  11 gtxrhsdd12(interior)
11 x 11 x  11 gtxrhsdd22(interior)
11 x 11 x  11 trkxrhs(interior)

CottonmouthZ4cxAnalysisGroup x z4cxconstraintsxgroup x group for function ‘z4cxconstraints‘. generated by einsteinengine. x

11 x 11 Language: 11 c
11 x 11 Type:  11 group

11

z4cxconstraintsxgroup x z4cxconstraints x function ‘z4cxconstraints‘ generated by einsteinengine. x

11 x 11 Language: 11 c
11 x 11 Reads:  11 atdd00(everywhere)
11 x 11 x  11 atdd01(everywhere)
11 x 11 x  11 atdd02(everywhere)
11 x 11 x  11 atdd11(everywhere)
11 x 11 x  11 atdd12(everywhere)
11 x 11 x  11 atdd22(everywhere)
11 x 11 x  11 theta(everywhere)
11 x 11 x  11 tmunubasex::ettt(interior)
11 x 11 x  11 tmunubasex::ettx(interior)
11 x 11 x  11 tmunubasex::etty(interior)
11 x 11 x  11 tmunubasex::ettz(interior)
11 x 11 x  11 tmunubasex::etxx(interior)
11 x 11 x  11 tmunubasex::etxy(interior)
11 x 11 x  11 tmunubasex::etxz(interior)
11 x 11 x  11 tmunubasex::etyy(interior)
11 x 11 x  11 tmunubasex::etyz(interior)
11 x 11 x  11 tmunubasex::etzz(interior)
11 x 11 x  11 chi(everywhere)
11 x 11 x  11 evoxgammatu0(everywhere)
11 x 11 x  11 evoxgammatu1(everywhere)
11 x 11 x  11 evoxgammatu2(everywhere)
11 x 11 x  11 evoxlapse(interior)
11 x 11 x  11 evoxshiftu0(interior)
11 x 11 x  11 evoxshiftu1(interior)
11 x 11 x  11 evoxshiftu2(interior)
11 x 11 x  11 gtdd00(everywhere)
11 x 11 x  11 gtdd01(everywhere)
11 x 11 x  11 gtdd02(everywhere)
11 x 11 x  11 gtdd11(everywhere)
11 x 11 x  11 gtdd12(everywhere)
11 x 11 x  11 gtdd22(everywhere)
11 x 11 x  11 trk(everywhere)
11 x 11 Type:  11 function
11 x 11 Writes: 11 hamcons(interior)
11 x 11 x  11 momconsu0(interior)
11 x 11 x  11 momconsu1(interior)
11 x 11 x  11 momconsu2(interior)
11 x 11 x  11 ztconsu0(interior)
11 x 11 x  11 ztconsu1(interior)
11 x 11 x  11 ztconsu2(interior)

11

CottonmouthZ4cxPostStepGroup x z4cxenforcexpt1xgroup x group for function ‘z4cxenforcexpt1‘. generated by einsteinengine. x

11 x 11 Language: 11 c
11 x 11 Type:  11 group

z4cxenforcexpt1xgroup x z4cxenforcexpt1 x function ‘z4cxenforcexpt1‘ generated by einsteinengine. x

11 x 11 Language: 11 c
11 x 11 Reads:  11 chi(interior)
11 x 11 x  11 evoxlapse(interior)
11 x 11 x  11 gtdd00(interior)
11 x 11 x  11 gtdd01(interior)
11 x 11 x  11 gtdd02(interior)
11 x 11 x  11 gtdd11(interior)
11 x 11 x  11 gtdd12(interior)
11 x 11 x  11 gtdd22(interior)
11 x 11 Type:  11 function
11 x 11 Writes: 11 chi(interior)
11 x 11 x  11 evoxlapse(interior)
11 x 11 x  11 gtdd00(interior)
11 x 11 x  11 gtdd01(interior)
11 x 11 x  11 gtdd02(interior)
11 x 11 x  11 gtdd11(interior)
11 x 11 x  11 gtdd12(interior)
11 x 11 x  11 gtdd22(interior)